40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Entry  Thu Sep 22 15:00:49 2016, Gabriele, Electronics, Configuration, Epics values now saved to frames 
    Reply  Fri Sep 23 08:33:39 2016, Gabriele, Electronics, Configuration, Epics values now saved to frames 
       Reply  Tue Sep 27 14:54:53 2016, Gabriele, Electronics, Configuration, Epics values now saved to frames 
Message ID: 129     Entry time: Tue Sep 27 14:54:53 2016     In reply to: 125
Author: Gabriele 
Type: Electronics 
Category: Configuration 
Subject: Epics values now saved to frames 

Apparently, there was a mismatch in the configuration, and DAQD was adding a wonderful 16 Hz comb all over the spectrum.

I stopped the processes, but couldn't restart x3cr1. It turned out that I can't save a channel to frames with a sampling frequency lower than 256 Hz. I changed the model, recompiled and restarted. Now the 16 Hz is gone.

Quote:

Now EPICS values are saved to frames, but they are all zero! I noticed that we always had the same problems with the cymac2 too.

So for the moment being I set up daqd to save X_NORM_IN1 and Y_NORM_IN1 at 32 Hz. In this way I can monitor the QPD centering.

Quote:

I was looking at some past trend data and discovered that EPICS values were not written to the frames. I added the following two lines to /opt/rtcds/tst/x3/target/fb/master to fix this:

/opt/rtcds/tst/x3/chans/daq/X3EDCU_CR1.ini
/opt/rtcds/tst/x3/chans/daq/X3EDCU_TST.ini

 

 

ELOG V3.1.3-